OpenStudy (anonymous):

Factor the trinomial: 7x^2 + 41x - 6

OpenStudy (jiteshmeghwal9):

7x^2+41x-6 => 7x^2+42x-x-6

OpenStudy (jiteshmeghwal9):

Now take 7x^2+42x. what is common in this equation ?

OpenStudy (anonymous):

I get the idea of finding numbers for m and n and have factored it out to 7x(x+6) - (x-6). I think that is right?

OpenStudy (jiteshmeghwal9):

yes u r right but u have done a mistake 7x(x+6)-1(x+6)

OpenStudy (jiteshmeghwal9):

got it ?

OpenStudy (anonymous):

7x^2 + 41x - 6 (x + 6)(7x - 1)

OpenStudy (anonymous):

That is where I'm lost .....how did you you get (x+6)(7x-1)

OpenStudy (anonymous):

do you know the foil method ?

OpenStudy (jiteshmeghwal9):

7x(x+6)-1(x+6) take (x+6) common (x+6)(7x-1)

OpenStudy (jiteshmeghwal9):

gt it ?

OpenStudy (anonymous):

jitesmeghwal explains better then I do

OpenStudy (jiteshmeghwal9):

Thanx buddy :)

OpenStudy (anonymous):

Ok...the Foil method is first, outside, inside, last

OpenStudy (anonymous):

only telling the truth :)

OpenStudy (anonymous):

yes...first, outside, inside, last

OpenStudy (anonymous):

Ok...(x+6)(7x-1) = 7x^2-x+42x-6 which in return is 7x^2+41x-6

OpenStudy (jiteshmeghwal9):

a(b+c)=a(b)+a(c)=ab+ac This is known as FOIL method :)

OpenStudy (anonymous):

yes...that is how I do it